# Modular Compilation for Quantum Chiplet Architectures

### v2

@@ 5.2. Benchmarks
Figure 5. Examples of the Benchmark Circuits
For our benchmark suite, we leverage a subset of circuits from the Supermarq (supermarq) suite. The circuits are chosen to be representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our benchmark suite comprises the following quantum circuits.
@@
Greenberger–Horne–Zeiling (GHZ). Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage. The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$. The algorithmic circuit structure is essentially a single chain of CNOT gates, as shown in Figure 5(c) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial. An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets. In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated, as shown in Figure 5(d). This results in a quantum circuit that is essentially two CNOT chains separated by single-qubit operations. The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing). Moreover, from a qubit routing perspective, the second CNOT chain forces the compiler to consider how qubit permutations in one circuit layer propagate to future layers.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials. The circuit for Hamiltonian Simulation, shown in Figure 5(e), resembles a slightly more complex GHZ circuit, including how it is completely serial. Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
@@ 5.3. Simulated Quantum Device Specifications
(backend_specs) $25$ $0.109\%$
SX (backend_specs) $25$ $0.109\%$
$\text{R}_{\text{z}}(\phi)$ (backend_specs) 00 $0.00\%$
CZ (intra-chiplet) (backend_specs) $34$ $0.605\%$
@@
Reset (backend_specs) $500$ $0.186\%$
Measure (backend_specs) $500$ $0.196\%$
(b) Instruction Properties.
@@
The topology of the simulated device is composed of a grid lattice of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex), as shown in Figure 6. The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations. Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
@@ 6. Experimental Results
Figures 79 present measurements of the recurring compilation time (i.e., solve time) of the baseline Strawman chiplet compiler, and SEQC’s two compilation stages: the non-recurring stratification stage and the recurring elaboration stage.
Figure 7. SEQC stratification time normalized to the 20-qubit (i.e., 2-chiplet) trial.
The average stratification time of SEQC across the benchmark suite obeys a quadratic trajectory ($R^{2}=0.9971$) with respect to the circuit size and number of chiplets, as shown in Figure 7. We note that the compilation time of the baseline, presented in Figure 8, also scales quadratically ($R^{2}=0.9964$). Although stratification time represents a one-time cost, the quadratic scaling of stratification points toward the practicality of SEQC. The stratification time depends heavily on the complexity of the input circuit. Simple circuits, such as GHZ or HamiltonianSimulation, exhibit a very gradual increase in stratification time as circuit size increases, while others, such as BitCode and PhaseCode, are much steeper. This observation holds true for the solve time of the baseline compiler as well. Collectively, these results indicate that the computational complexity of the baseline compiler and the stratification stage of SEQC are affected similarly by the same quantum circuit characteristics.
Figure 8. Baseline compilation time normalized to the 20-qubit (i.e., 2-chiplet) trial.
Figure 9. Relative elaboration time of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
From Figure 9, we observe an improved SEQC elaboration time compared to the baseline solve time, leading to a roughly 24×2-4$\times$2 - 4 × speedup, up until $800$-qubit (i.e., $80$-chiplet) trials. The observed speedup reflects the expected performance improvement that our stratify-elaborate technique should garner from parallelization and problem-space reduction. For $81$-chiplet experiments and above, our results begin to reflect a far more marginal speedup compared to baseline, reflecting the limitations of our evaluation hardware. Namely, as the number of chiplets begins to approach the number of physical cores and hardware threads on our hardware, SEQC experiences difficulties in evenly distributing its tasks. For example, under-allocating tasks causes poor core utilization, while over-allocation leads to process thrashing, either way resulting in poorer elaboration time. However, in all cases, SEQC outperforms the baseline on average.
On a circuit-by-circuit basis, we notice higher relative performance with parallel workloads (like BitCode and PhaseCode) compared to serial workloads (GHZ and HamiltonianSimulation). Notably, above the $81$-chiplet threshold where the classical machine we use for compilation no longer suffices, elaboration time exceeds the baseline solve time for the serial workloads. This is because serial circuits are fundamentally trivial to solve, and thus benefit less from the parallelism offered by SEQC. We verified experimentally that in these cases, SEQC spends almost all of its time on thread startup and teardown, rather than useful work. Moreover, the raw compilation time is trivially small in these cases, owning to the simplicity of the circuits, thus relative performance differences have little practical significance.
Figure 10. Relative estimated fidelity (ESP) of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (higher is better).
Figure 10 presents the relative estimated fidelity of SEQC compared to the baseline with respect to the number of qubits per circuit. Notably, as the number of chiplets / the circuit size increases, SEQC improves its relative fidelity against the baseline. We begin to see higher fidelity circuits in the geometric mean once the circuit size exceeds $36$ chiplets, up to a maximum of $36\%$ higher fidelity. All benchmarks reflect higher relative performance for the $50$-chiplet trials and above.
Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 11 presents the estimated per-shot execution time of the circuits produces by SEQC compared to the circuits produced by the baseline compiler, as a function of the number of qubits per circuit. As explained above, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets). Under this model, we observe execution time benefits up to a 1.92×1.92$\times$1.92 × improvement over baseline at $750$ qubits ($75$ chiplets). Not only are improvements in execution time desirable for quantum circuit throughput, they could also result in lower decoherence and thus improved circuit fidelity.
@@
Figure 12. Relative number of inter-chiplet gates for SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
Figure 12 shows the number of inter-chiplet swaps in the circuits generated by SEQC relative to the circuits generated by the baseline, with respect to the number of qubits per circuit. For essentially all tested circuits, SEQC achieves a significant reduction in the number of inter-chiplet swap gates, appearing to asymptotically converge on a roughly 4×4$\times$4 × reduction, with the greatest reduction of 4.6×4.6$\times$4.6 × at $800$ qubits / $80$ chiplets. This aligns with our model of inter- vs intra-chiplet gates, where inter-chiplet SWAPs are assumed to have 4×4$\times$4 × the error (smith_2022_chiplets) and 4×4$\times$4 × the gate duration.
Figure 13. Relative circuit depth of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
@@
Figure 14. Relative gate count of SEQC normalized to the Strawman baseline for devices with 10-qubit chiplets (lower is better).
Relative circuit depth and gate count, shown in Figures 13 and 14, respectively, share similar behavior and analysis. At small circuit sizes, SEQC has higher circuit depth and gate count than the baseline. The gap shrinks as the circuit size increases, although SEQC always tends to perform worse on these metrics on average. This behavior is to be expected, since the baseline directly optimizes for circuit depth and gate count, while SEQC prioritizes other metrics, e.g., reducing inter-chiplet gates.
6.1. Discussion
The discrepancy between the circuit depth and gate count results compared to the execution time and fidelity results, respectively, are indicative of the differences between SEQC and the contemporary quantum compilers. Circuit depth and gate count represent the heuristics used in the baseline, especially SABRE (sabre_swap), and they are fundamentally unweighted metrics. In other words, circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. Meanwhile, SEQC is able to more directly estimate execution time and fidelity, the practically relevant metrics, by constructing weighted heuristics.
With the context of our static metrics (number of inter-chiplet gates, circuit depth, and gate count), it is possible to justify the improvements in relative fidelity and relative per-shot execution time as the circuit size increases. Namely, as expected, the unweighted heuristics in SABRE tend to under-represent the error and gate delay contribution of inter-chiplet gates. This effect gets exacerbated for wider or deeper circuits, which increase the number of opportunities for cross-chiplet interactions.
@@ 6.1. Discussion
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increase, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates. These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
@@ 7. Related Work
Even though quantum computing is still nascent, there is a significant body of work on techniques for compiling quantum circuits for NISQ processors. Here, we focus on the work that best relates to our contributions.
MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software. By entangling ancillary qubits and distributing them throughout the chiplet architecture’s topology, the work virtualizes those qubits into fast network-on-chip routing channels without hardware modifications. Future work to extend SEQC with MECH’s methodology may enable combining MECH’s techniques with SEQC’s allocation procedures to gain the benefits of each.
@@
Noting the scaling challenges and fidelity issues associated with large, monolithic quantum machines, CutQC (cutqc) establishes a hybrid computing paradigm which enables smaller-scale quantum computers to evaluate quantum circuits they would have otherwise had too few qubits to execute. To establish said paradigm, the work introduces a novel circuit-cutting technique for partitioning larger quantum circuits into small subcircuits. Classical post-processing is performed on the results of the subcircuit evaluations, merging them to find the result that would be generated by the uncut circuit. Circuit cutting as performed in CutQC served as an inspiration for SEQC’s stratification step, but unlike CutQC, SEQC does not perform any classical “stitching” of the results afterwards. It is interesting to note that our first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
Bandic et al.(bandic2024profiling) also note the benefits of modularity. Their study is motivated by application-specific optimization – they seek techniques that allow a quantum circuit’s structure to influence hardware architecture. They define the relationship between circuit structure and mapping efficiency, utilizing this information to optimize the architecture of single- and multi-core quantum processors.
Piveteau and Sutter (circuit_knitting) investigate the benefits of using quasiprobability simulation to allow subcircuits cut from larger quantum circuits to perform their quantum operations independently, simulating inter-chiplet link behavior where the subcircuits connect. When augmented with classical communication between circuits to reduce the simulation overhead incurred by sampling, they demonstrate that resulting circuits reach fixed accuracy with less simulation time. Unlike SEQC that targets quantum compilation, this technique has been developed to partition large quantum circuits into subcircuits that fit on smaller devices, at the cost of a simulation overhead to classically combine them.
Lin et al. (let_every_qubit) recognize that each 2-qubit gate in a quantum circuit has a dynamic behavioral profile which depends on manufacturing and environmental factors, and present a compilation technique which uses each 2-qubit set’s calibration data to determine potentially nonstandard 2Q gate operations which are much more ideal for that set to perform than their standard counterparts—essentially, choosing a more favorable basis. By characterizing these nonstandard gates, using their unitaries to synthesize standard gates, and essentially allowing 2Q gates to perform standard operations along bases more compatible with their respective physical states, the technique improves fidelity. This technique is orthogonal to SEQC and could work synergistically with it.
@@
By mapping circuit partitioning to a minimum cut problem, Bandic et al. (bandic2023mapping) use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between circuit partitions. QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$. SEQC, in contrast, does not require acceleration through quantum means, but at the cost of not guaranteeing an optimal solution (it is rather a “best-effort” approach that works well in practice).
Through discussion of scalability issues impeding the development of larger monolithic quantum circuits, Ovide et al. (ovide2023mapping) motivate the necessity of mapping quantum circuits to multi-core architectures and present a commonly agreed-upon likely computing paradigm which leverages both quantum and classical methods to foster inter-chip communication, circumventing the scalability challenges inherent to monolithic quantum processors. While this work advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and then routing unsolved. SEQC provides an algorithm and implementation, addressing both of these challenges.
Using iterative optimization via the Hungarian algorithm, Hungarian Qubit Assignment (HQA) (escofet2023hungarian) seeks to optimize the qubit-to-core mapping problem. Because the Hungarian algorithm is here extended with the ability to adjust a qubit’s cost matrix according to behavior multiple timeslices ahead, the model can leverage lookahead to optimize further the number of inter-core swaps. However, HQA operates under the assumption that all inter-chiplet communications have the same cost, meaning chiplets must be fully connected. SEQC, on the other hand, is able to work with any arrangement of chiplets and inter-chiplet links so long there is a path between any two chiplets, taking into account the distance between pairs of chiplets to generate improved circuits.
Iterating on their previous work on Hungarian Qubit Assignment, Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds. This is significant, since it enables optimality assessment for solutions to the qubit mapping problem.
Pastor et al. (pastor2024circuit) leverage deep reinforcement learning to partition quantum circuits for execution across multiple processing cores. Using a DRL model enhanced with policy-masking mechanisms and a lookahead-capable observation array, the work finds valid qubit-circuit mappings for each of timeslice while prioritizing the reduction of inter-core swaps. Notably, it only considers inter-core swaps for optimization, deeming intra-core single-and-multiple qubit operations so inexpensive that their optimization should serve as fodder for inter-core swap reductions. This is sensible, but leaves optimization potential on the table, since it only seeks to generate valid mappings within each timeslice. Additionally, similar to HQA, this technique treats all inter-chiplet communications equally, further missing optimization opportunities with non-fully connected chiplet topologies. SEQC, in contract, is cognizant of the differing costs of moving information between each pair of chiplets, and optimizes accordingly.
To solve the qubit allocation (mapping) problem, Liu et al. (liu_2023_tackling) develop an algorithm that combines partitioning, permutation-aware synthesis, and a modified version of SABRE. In a bottom-up approach, this technique partitions a logical circuit into blocks, performs permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally uses an augmented SABRE algorithm to perform initial layout and routing between blocks. In contrast, SEQC takes a top-down approach, where SWAP permutations between chiplets are settled prior to any compilation tasks performed within each block. As a consequence, we only need to generate solutions for one qubit layout per block, rather than synthesizing all possible block permutations.
@@
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike qubit swap optimization, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.

### v3

@@ 4.1. Benchmark Suite
Figure 5. Examples of benchmark applications for 5-qubit circuits.
We leverage circuits from Supermarq (supermarq) that are representative of practical, real-world applications for quantum computing, while also being feasible to compile for a large number of qubits. In particular, our suite comprises the following quantum circuits.
@@
Greenberger–Horne–Zeiling (GHZ). Entanglement is one of the fundamental properties of quantum states and a major source of quantum advantage. The GHZ state represents a maximally entangled $n$-qubit state of $\frac{1}{\sqrt{2}}\left(\ket{|}0>^{\otimes n}+\ket{|}1>^{\otimes n}\right)$. The algorithmic circuit structure is essentially a single chain of CNOT gates (Figure 5(c)) From a compilation perspective, this represents a pure exercise in qubit allocation, because the logical circuit is entirely serial. An $n$-qubit GHZ circuit that fully utilizes a $k$-chiplet device should necessitate $\lfloor\frac{n}{k}\rfloor-1$ logical CNOT operations to extend between chiplets. In turn, this will necessitate at least $2\left(\lfloor\frac{n}{k}\rfloor-1\right)$ inter-chiplet SWAPs to transfer qubit in and out of a chiplet.
VQE. The variational quantum eigensolver (VQE) (vqe) is another hybrid quantum-classical variational algorithm that is often used in chemistry applications. In our experiments, we choose a $2$-layer ansatz whose parameters are again randomly generated (Figure 5(d)). The resultant circuit structure features two CNOT chains separated by single-qubit operations. The CNOT chains can be interleaved, which in conjunction with the single-qubit gates, creates many opportunities for parallelism (resembling pipelining in classical computing). Moreover, from a qubit routing perspective, the second CNOT chain forces the compiler to consider how qubit permutations in one circuit layer propagate to future layers.
Hamiltonian Simulation. Quantum devices are naturally well-suited to simulating Hamiltonians. The Supermarq benchmark selects the 1D Transverse Field Ising Model (TFIM) system to model, which is relevant for phase transitions in magnetic materials. The circuit for Hamiltonian Simulation (Figure 5(e)) resembles a slightly more complex GHZ circuit, including how it is completely serial. Compared to GHZ, this circuit should test the compiler’s ability to lookahead during routing to detect back-to-back gates on the same two-qubit pair, while also emphasizing parallelism originating from the efficient placement of SWAP gates.
@@ 4.2. Simulated Quantum Device Specifications
(backend_specs) $25.0$ $0.109\%$
SX (backend_specs) $25.0$ $0.109\%$
$\text{R}_{\text{z}}(\phi)$ (backend_specs) $0.0$ $0.000\%$
CZ (intra-chiplet) (backend_specs) $34.0$ $0.605\%$
CZ (inter-chiplet) $126.0$ $2.420\%$
@@
Reset (backend_specs) $500.0$ $0.186\%$
Measure (backend_specs) $500.0$ $0.196\%$
(b) Instruction Properties
Universal gate set only. Non-universal gate set only.
@@
The simulated device has a grid lattice topology of symmetric chiplet modules, where each chiplet features a heavy-hexagon qubit topology (heavy_hex) (Figure 6). The intra-chiplet topology is constructed such that the heavy-hex structure is preserved for the global device topology (ignoring chiplet division). The tessellation requirements of each chiplet restrict valid chiplet sizes to multiples of $10$; however, for our experiments, we limit the size of each chiplet to the minimum of $10$ qubits due to hardware limitations. Finally, note that for each valid chiplet count, there may exist multiple valid chiplet topologies. In this work, we generate our backends to use the “most-square” topology for a given chiplet count. For example, a $12$-chiplet machine would be generated with a grid of $3\times 4$ chiplets.
@@ 5. Experimental Results
We present results for the Qiskit reference compiler, the peephole-augmented Qiskit (PA-Qiskit) compiler, the SEQC compiler using simulated annealing (SEQC-AN), and the SEQC compiler using STA (SEQC-STA). The Qiskit-based and SEQC compilers prioritize different heuristics in compilation. We examine how well each compiler is able to minimize its preferred heuristics, and how well that translates to practical solution quality. Finally, we analyze compilation time, specifically the recurring and non-recurring compilation times for repeating compilations of a single application.
Number of inter-chiplet gates
@@
Figure 7. Relative number of inter-chiplet gates for SEQC normalized to the Qiskit baseline (lower is better).
The primary heuristic we minimize in SEQC is the number of inter-chiplet gates in a quantum circuit. This is because we expect that the inter-chiplet gates of a chiplet architecture contribute the most to error and execution time relative to other gates (at least by a factor of 4×4$\times$4 × (smith_2022_chiplets)). As a result, the number of inter-chiplet gates should represent a strong proxy metric for the quality of a compiler solution, and thus minimizing the number of inter-chiplet gates should result in higher-fidelity, faster-executing quantum circuits. Accordingly, SEQC consistently produces results that have significantly fewer inter-chiplet gates compared to Qiskit-based compilers (Figure 7). Specifically, compilation with SEQC-STA results in $82.43\%$ fewer inter-chiplet gates on average compared to Qiskit (up to $92.94\%$). Compared to PA-Qiskit, compilation with SEQC-STA results in $64.92\%$ fewer inter-chiplet gates on average, up to $78.62\%$. It is worth mentioning that the inter-chiplet gates in the universal backend would be CZ gates, while the non-universal backend would be SWAP gates. For every inter-chiplet (CZ) gate in Qiskit, PA-Qiskit should have 2 inter-chiplet (SWAP) gates. However, we observe PA-Qiskit produces circuits with a roughly comparable number of inter-chiplet gates to Qiskit. In fact, as size increases, PA-Qiskit tends to result in fewer inter-chiplet gates than Qiskit. This outcome may indicate the presence of optimizations that are removing extraneous SWAPs.
Gate count and circuit depth.
@@
Figure 8. Relative gate count of SEQC normalized to the Qiskit baseline for devices (lower is better).
Figure 9. Relative circuit depth of SEQC normalized to the Qiskit baseline for devices (lower is better).
Gate count and circuit depth are static, backend-agnostic metrics that can serve as indicators for a quantum circuit’s expected fidelity and/or execution time, where circuit depth is the critical path of a quantum circuit in terms of individual gates (mike_and_ike). SABRE (zou2024lightsabrelightweightenhancedsabre; sabre_swap), moreover, explicitly minimizes gate count by incorporating it into its cost functions. We observe that SEQC typically produces results with higher gate count and circuit depth compared to Qiskit and PA-Qiskit (Figures 8 and 9). This is due to SEQC prioritizing the reduction of the number of inter-chiplet gates over total gate count or circuit depth. Nevertheless, SEQC-STA, in particular, is able to demonstrate competitive results in these metrics for larger circuits. In terms of relative circuit depth, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $500$ qubits (50 chiplets) and above. In terms of relative gate count, SEQC-STA has comparable or slightly better results compared to PA-Qiskit at $400$ qubits (40 chiplets) and above. For the $810$ and $1000$ qubit ($81$ and $100$ chiplet) trials, it achieves $14.11\%$ and $8.87\%$ gate count reduction over Qiskit, respectively, despite the latter using universal inter-chiplet links.
Fidelity
Figure 10. Relative estimated fidelity (ESP) of our proposed compilers normalized to the Qiskit baseline (higher is better).
Practical applications of quantum computing require some guarantee of program output correctness (measurements). Thus, the highest priority for quantum circuit compilation is to maximize fidelity. Fidelity is impacted by the decoherence constraints of individual qubits, as well as the errors on a gate-by-gate basis (mike_and_ike). Due to hardware limitations, we were unable to estimate fidelity though simulation. Instead, we leverage estimated success probability (ESP) (esp) to approximate fidelity, which can be computed statically. Specifically, we track the per-qubit ESP of a quantum circuit and report an averaged ESP for each benchmark from the geometric mean of the circuit measurements. Figure 10 presents the relative estimated fidelity of each of our compilers compared to the Qiskit compiler with respect to circuit size. There is a noticeable benefit in fidelity associated with the SEQC compiler, particularly with SEQC-STA. We observe an average improvement of $22.13\%$ over PA-Qiskit, up to a $62.87\%$ improvement. Moreover, it is important to mention that the Qiskit trial was ran on a backend with universal inter-chiplet links, which should translate to higher fidelity solutions. Nevertheless, SEQC-STA demonstrates an average improvement of $9.30\%$, up to a $49.99\%$ improvement.
Estimated execution time.
Figure 11. Estimated per-shot execution time of circuits produced by SEQC normalized to circuits produced by the Qiskit baseline (lower is better).
The per-shot execution time of a quantum circuit is determined by the critical path of gate operations weighted by gate duration. In addition to performance concerns, execution time indirectly affects fidelity on NISQ-era devices due to decoherence. In other words, it is imperative that quantum circuits execute within the decoherence time to ensure high-fidelity outcomes. Figure 11 presents the estimated per-shot execution time of our proposed compilers relative to the reference as a function of the number of qubits per circuit. As explained previously, we approximate the duration of an inter-chiplet SWAP as 4×4$\times$4 × the duration of an intra-chiplet SWAP (smith_2022_chiplets). Under this model, we observe that the SEQC compilers consistently outperform PA-Qiskit in execution by $24.19\%$ on average ($63.98\%$ max). It is notable that, despite the gate set affordances given to the Qiskit compiler, SEQC-STA is able to achieve parity in execution time starting from $600$ qubits ($60$ chiplets). For $810$ qubits ($81$ chiplets) and above, SEQC-STA actually compiles faster-executing quantum circuits.
Compilation time
Figure 12. Compilation time of our evaluated compilers (lower is better).
Figure 13. Relative recurring compilation time of our proposed compilers normalized to the Qiskit baseline (lower is better).
Near-term quantum devices feature rapidly fluctuating error characteristics that can change hourly, or even on a run-to-run basis (Tannu_2019_ErrorVariability; Murali_2019_NoiseMapping; dasgupta2021stability). Thus, across the entire execution lifetime of an application (quantum circuit), regular re-compilation is necessary to generate the highest-quality circuits with the most up-to-date characteristics possible for every execution. Traditional quantum compilers, like Qiskit and PA-Qiskit, approach each compilation tabula rasa, where each new compilation restarts from the original algorithmic circuit. It is challenging to speed up repeated recompilations without fundamental changes to their design, as noise-aware routing and synthesis algorithms, which leverage updated error characteristics, usually function from a device-agnostic starting point. In contrast, SEQC’s stratification stage enables stateful retention of previous compilations. It is agnostic to the ever-changing error characteristics of a quantum device, and instead only depends on the device’s static topology. Therefore, a stratified quantum circuit can be perennially reused over multiple rounds of compilation, and thus SEQC only repeats the elaboration stage.
Nevertheless, although stratification is a one-time cost for the SEQC compiler, it must be scalable in order to be practical. In Figure 12, we plot the non-recurring compilation time of our evaluated compilers. For our SEQC compilers, we plot stratification time and stratification + elaboration time. These can be compared to the total solve time for the Qiskit and PA-Qiskit compilers. We demonstrate that all compilation times obey a quadratic trajectory ($R^{2}\geq 0.9928$).
Figure 13 shows the recurring compilation time. Specifically, this refers to the elaboration time of the SEQC compilers and the full compilation time for Qiskit and PA-Qiskit. Overall, SEQC compilers have a large speedup over Qiskit-based compilers, owing to the use of parallelism and modular compilation strategies. Compared to Qiskit, SEQC-STA achieves 3.27×3.27$\times$3.27 × speedup on average (up to 6.74×6.74$\times$6.74 ×), and compared to PA-Qiskit 4.69×4.69$\times$4.69 × average speedup (6.61×6.61$\times$6.61 × max).
6. Discussion
These results demonstrate a flaw with gate count and circuit depth as heuristics: their lack of noise-awareness. Circuit depth and gate count only consider the number of gates (in the critical path and in total) irrespective of each one’s unique underlying properties. It is expected that two-qubit gates will be significantly more noisy than single-qubit gates, and inter-chiplet gates should be significantly more noisy than intra-chiplet gates (smith_2022_chiplets). Failure to account for duration and error differences between gates or types of gates in quantum compilation heuristics will inevitably result in sub-optimal solutions when the discrepancies are extreme, e.g., in chiplet architectures. Conversely, the hierarchical structure of the SEQC framework with the stratification and elaboration stages imparts some degree of implicit noise-awareness, where minimizing the number of inter-chiplet gates takes precedence over intra-chiplet/total gates. As a result, SEQC can better optimize for the practically significant metrics of circuit fidelity and circuit execution time.
@@ 6. Discussion
Overall, we observe that SEQC yields circuits with more desirable characteristics when executing on modular quantum architectures, while at the same time it achieves faster compilation times. Also, as the number of qubits increases, the performance improvement of SEQC widens compared to the baseline, especially for the most important metrics of fidelity, compiled circuit execution time, and number of inter-chiplet gates. These trends hold promise in allowing for faster and better compilation for future large-scale modular quantum devices than we can achieve with today’s frameworks.
@@ 7. Related Work
MECH (mech) sacrifices some ancillary qubit resources to build an ”adjustable multi-entry communication highway” through software. By entangling ancillary qubits and distributing them throughout a chiplet, MECH virtualizes those qubits into fast network-on-chip routing channels. Lin et al. (let_every_qubit) present a compilation technique that utilizes calibration data to determine potentially nonstandard 2Q gates that can perform standard operations along bases more compatible with their respective physical states and improve fidelity. Bandic et al.(bandic2024profiling) utilize the relationship between quantum circuit structure and mapping efficiency to optimize the architecture of monolithic and modular quantum processors, on the path to application-specific hardware optimization. Escofet et al. (escofet2024revisiting) develop a characterization technique for arbitrary quantum circuits which yields theoretical upper and lower optimization bounds, enabling the optimality assessment of qubit mapping solutions. All these techniques are orthogonal to SEQC, and could work synergistically with it to achieve the combined benefits.
@@
CutQC (cutqc) enables smaller-scale quantum computers to execute quantum circuits that are otherwise too large to fit on the device. It partitions large quantum circuits into smaller subcircuits, executes them one at a time, and then merge the individual results at the cost of classical post-processing. Piveteau and Sutter (circuit_knitting) similarly allow subcircuits to execute independently, and use quasiprobability simulation to simulate inter-chiplet link behavior where the subcircuits connect, reaching fixed accuracy with less simulation time. SEQC’s stratification stage was inspired by CutQC, but unlike these works SEQC does not perform any classical “stitching” or simulation. It is interesting to note that the first version of SEQC was formulating the problem using constrained optimization as well, but the slowdown of the stratification step was unacceptable and impractical for large quantum circuits, and thus was abandoned in favor of the heuristics we currently employ in our compiler.
@@
Bandic et al. (bandic2023mapping) map circuit partitioning to min-cut and use the QUBO model with quantum solvers to generate a set of cuts intended to minimize swaps between partitions. QUBO employs quantum optimization algorithms on quantum annealers to accelerate the solution process and remain practical, as the search space for transmission qubits is $2^{n}$. Instead, SEQC takes a “best-effort” approach that works well in practice, and does not require acceleration through quantum means.
Some works on qubit mapping share a few characteristics with SEQC. In a bottom-up approach to qubit mapping, Liu et al. (liu_2023_tackling) partition a circuit into blocks that span an identical number of qubits, perform permutation-aware synthesis on each block in parallel (solving for all possible qubit topologies), and finally use an augmented SABRE algorithm to map the blocks to the target device and route between blocks. SEQC takes a top-down approach instead, where SWAP permutations between chiplets are settled prior to any compilation tasks performed for each chiplet (block). As a result, SEQC only needs to solve for one qubit layout per block, rather than all permutations. Hungarian Qubit Assignment (HQA) (escofet2023hungarian) extends the Hungarian algorithm with the ability to adjust a qubit’s cost matrix according to behavior multiple time slices ahead. Pastor et al. (pastor2024circuit) use deep reinforcement learning to partition quantum circuits for execution across multiple chiplets. Both works aim to reduce inter-chiplet swaps, but assume that all inter-chiplet communications have the same cost, i.e., chiplets are fully connected, which is both impractical for large-scale systems and misses optimization opportunities. SEQC, in contrast, can work with arbitrary chiplet topologies. Ovide et al. (ovide2023mapping) leverage both quantum and classical methods for inter-chiplet communication to address the scalability challenges of monolithic processors. While it advocates for a two-level, hierarchical approach for quantum circuit mapping, it leaves the actual problem of qubit placement and routing unsolved. SEQC provides an algorithm and implementation, addressing both challenges.
@@
Parallelizing SEQC’s recurrent elaboration step was inspired by traditional classical compilation techniques, which has a decades-long history and is now commonplace (Mickunas_1978_ParallelCompilation; Gross_1989_ParallelCompilation). Besides traditional compilation, though, we can also draw parallels between our work and prior research on data movement in the classical space. Data movement is widely understood to be the primary performance bottleneck in the classical architecture space. In the last two decades, work has proliferated on dataflow and spatial architectures, much of which focuses on mapping problems. Consider neural network architectures, for example, which seek to map their physical topologies as closely as possible to problem topologies. Eyeriss (Eyeriss_2016_Chen) approaches the dataflow mapping problem by proposing a row-stationary architecture, and then optimizing the mapping of problem to processing element such that elements which are spatially local in the problem are spatially local in hardware during execution. This is not unlike the target optimization for the stratification stage, and in fact is performed for many of the same reasons — like inter-chiplet swaps, data movements between more distant classical programming elements in a dataflow architecture are more expensive, and minimizing their frequency yields performance improvements.